Consider the following income statement for the Heir Jordan Corporation: HEIR JORDAN CORPORATION Income Statement Sales $ 46,500 Costs 36,100 Taxable income $ 10,400 Taxes (30%) 3,120 Net income $ 7,280 Dividends $ 3,100 Addition to retained earnings 4,180 The balance sheet for the Heir Jordan Corporation follows. HEIR JORDAN CORPORATION Balance Sheet Assets Liabilities and Owners’ Equity Current assets Current liabilities Cash $ 2,600 Accounts payable $ 2,400 Accounts receivable 3,700 Notes payable 5,200 Inventory 9,000 Total $ 7,600 Total $ 15,300 Long-term debt $ 24,000 Owners’ equity Fixed assets Common stock and paid-in surplus $ 17,000 Net plant and equipment $ 38,200 Retained earnings 4,900 Total $ 21,900 Total assets $ 53,500 Total liabilities and owners’ equity $ 53,500 Prepare a pro forma balance sheet, assuming a 10 percent increase in sales, no new external debt or equity financing, and a constant payout ratio. (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.) HEIR JORDAN CORPORATION Pro Forma Balance Sheet Assets Liabilities and Owners’ Equity Current assets Current liabilities Cash $ Accounts payable $ Accounts receivable Notes payable Inventory Total $ Total $ Long-term debt $ Owners’ equity Fixed assets Common stock and paid-in surplus Net plant and equipment $ Retained earnings Total $ Total assets $ Total liabilities and owners’ equity $ Calculate the EFN. (Negative amount should be indicated by a minus sign. Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.) EFN $
Answer #1
Heir Jordan Corporation
Dividend in percentage of Net Income =($3100/$7280)*100 43%
Retained Earnings in percentage of Income=($4180/$7280)*100 57%
Current Performa Income Statement
Sales(A) $         46,500.00 46500*1.10 $                    51,150
Operating Cost(B) $         36,100.00 36100*1.10 $                    39,710
EBT(D )=(C)=(A)-(B) $         10,400.00 $                    11,440
Taxable Income=(D)=(C )*30% $           3,120.00 $                      3,432
EAT=(E )=(C )-(D) $           7,280.00 $                      8,008
Dividend=(I)=(E)*43% $           3,100.00 8008*43% $                      3,443
Retained Earnings=(H)-(I) $           4,180.00 $                      4,565
Current Balance Sheet Performa Balance Sheet
Assets
Cash $           2,600.00 2600*1.10 $                2,860.00
Accounts Receivable $           3,700.00 3700*1.10 $                4,070.00
Inventories $           9,000.00 9000*1.10 $                9,900.00
Net Fixed Assets $         38,200.00 38200*1.10 $              42,020.00
Total Assets=(A) $         53,500.00 $                    58,850
Liabilties
Accounts Payable $           2,400.00 2400*1.10 $                      2,640
Notes Payables $           5,200.00 $                      5,200
Long Term debt $         24,000.00 $                    24,000
Common Stock $         17,000.00 $                    17,000
Retained Earnings=($4900+$4565) $           4,900.00 $                      9,465
Total Liabilities & Equity=(B) $         53,500.00 $                       -   $                    58,305
AFN=(A)-(B) $                          545
AFN $                     545
If 10% increase in sales Note Payable will be($5200+$545) $                 5,745